OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the function y = –x² + 6x –11? (3, –2) (3, 16) (–3, –29) (–3, –20)

OpenStudy (accessdenied):

same equation applies here for the x-value of the vertex x = -b/2a we can then re-insert that value of x into our function to find the corresponding y-value (-b/2a, f(-b/2a))

OpenStudy (accessdenied):

so in this case... x= -(6) / 2(-1) = -6/(-2) = 3 -(3)^2 + 6(3) - 11 -9 + 18 - 11 9 - 11 -2 The point would be (3, -2)